OpenStudy (fanduekisses):

55 is 10% of how much?

OpenStudy (fanduekisses):

How do I find the answer?

OpenStudy (anonymous):

I think its 550

OpenStudy (anonymous):

how much do you have to multiply 10% by in order to get 100%?

OpenStudy (anonymous):

55 divided by 10% on a calculator

OpenStudy (anonymous):

teach, don't give

OpenStudy (fanduekisses):

Thanks guys ;)

OpenStudy (anonymous):

55 = 10% *10 *10

OpenStudy (fanduekisses):

So I have to multiply 10 times?

OpenStudy (shamil98):

0.10 x = 55 x = 55/0.10

OpenStudy (anonymous):

10% * 10 =100% 55 * 10 =

OpenStudy (anonymous):

let total be x 10% of x = 55 10/100 * x = 55 1/10 * x = 55 x = 55 * 10 = 550

OpenStudy (abb0t):

9,000,000,000,000,000,000,001

OpenStudy (anonymous):

9,000,000,000,000,000,000,001.012* @abb0t